Can some explain what the correct answer is? They are the ones with the red x. Thanks
Award: 5 out of 10.00 points You are considering a stock investment in one of two firms (NoEquity, Inc., and NoDebt, Inc.), both of which operate in the same industry and have identical operating income of $32.5 million. NoEquity, Inc., finances its $65 million in assets with $64 million in debt (on which it pays 10 percent interest annually) and $1 million in equity. NoDebt, Inc., finances its $65 million in assets with no debt and $65 million in equity. Both firms pay a tax rate of 30 percent on their taxable income. Calculate the net income and return on assets for the two firms. (Enter your dollar answers in millions of dollars. Round all answers to 2 decimal places.) No Equity S 21.65 m No Debt $22.25m Net income Return on assets 33.30 % 35.00 % eBook &Resources
Answer #1

Operating Income = $32.5 Million; firm pays a tax rate of 30 percent on their taxable income

Net income for equity = Operating income - Taxes (30% tax on operating income)

Net income for equity = Operating income - Taxes (30% tax on operating income) - 10% interest on debt

NoEquity - $65 Million in finances

  • $64 Million in debt, $1 Million in equity; 10 percent interest annually
  • Net income - $32.5 - $9.75 Million - 100000 = $21.75 - 100000 = $21.65 Million
  • Return on assets - 21.65 /65 = 33.3%

NoDebt - $65 Million in finances

Net Income = $32.5 - $6.4 = $26.1Million (Income - Interest on $64 million debt)

$26.1- $7.8 Million = $18.3 Million (Deducting 30% tax on $26.1 Million)

Return on assets - 18.3/65 = 28.15%